A rock is thrown horizontally from the top of a cliff 88 m high, with a horizontal speed of 25 m/s. For what interval of time is the rock in the air


1
Expert's answer
2021-04-29T10:40:59-0400

"s=ut+ \frac{1}{2}at^2"

"88=0t+ \frac{1}{2} \times 9.8t^2"

"88= \frac{1}{2} \times 9.8t^2"

"88 \times 2 \times \frac{1}{9.8}= t^2"

"t=\sqrt{\frac{88 \times 2}{9.8}} =4.24 s"